a train left a station traviling 100km/hr. a second train left 2 hr later and headed in the same direction at 125km/hr. After how many hours did the second train overtake the first train?
Answer by checkley75(3666)

100T=125(T-2)
100T=125T-250
125T-100T=250
25T=250
T=250/25
T=10 IN 10 HOURS THEY WILL MEET.